OpenStudy (anonymous):

Please tell me what are the restrictions to this multiplication of ration expressions.

OpenStudy (anonymous):

\[\frac{ z^2 }{ z+1 } * \frac{ z^2 + 3z + 2 }{ z^2 + 3z }\]

OpenStudy (anonymous):

@satellite73 Is 0 one o the restrictions for z?

OpenStudy (anonymous):

I know that z cannot equal to -1 and -3. Can z equal to 0 or not?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

I know the result is this \[\frac{ z^2 + 2z }{ z + 3 }\]

jimthompson5910 (jim_thompson5910):

factor z^2 + 3z to get z(z+3) you'll find that z = 0 or z = -3 makes z^2 + 3z = z(z+3) equal to zero so yes, z = 0 is one of the restrictions

OpenStudy (anonymous):

So because if z = 0, the whole expression equaling to 0, makes 0 a restriction, right?

OpenStudy (anonymous):

How do I factor \[x^2 + 4x - 5\]

jimthompson5910 (jim_thompson5910):

find two numbers that multiply to -5 and add to 4

jimthompson5910 (jim_thompson5910):

list all the ways to multiply to get -5 -1*5 1*(-5) then see which pair of factors add to +4

OpenStudy (anonymous):

\[x(x+4) - 5\]

jimthompson5910 (jim_thompson5910):

which pair of factors listed above add to 4?

OpenStudy (anonymous):

1 and -5

jimthompson5910 (jim_thompson5910):

1 plus -5 = -4

OpenStudy (anonymous):

\[(x + 1)(x -5)\]

OpenStudy (anonymous):

yeah i meant that

jimthompson5910 (jim_thompson5910):

but we want the two numbers to add to +4

jimthompson5910 (jim_thompson5910):